OpenStudy (anonymous):

vdv/dx = 2(16+v^2) differential equation

OpenStudy (anonymous):

it is a fully seperatable DE so you should put all the v terms on one side and all the x terms on the other

OpenStudy (anonymous):

that's what i cant do, because there's v on both sides...

OpenStudy (anonymous):

do i divide by 2(16+v^2)?

OpenStudy (anonymous):

yes. \[\frac{ v dv }{ (16 + v^{2} ) } = 2 dx\]

OpenStudy (anonymous):

but how do i integrate the left hand side?

OpenStudy (anonymous):

is it 2vln(16+v^2)

OpenStudy (anonymous):

let u = 16 + v^2 --> du = 2v dv = \[1/2 \int\limits_{}^{} \frac{ du }{ u }\]

OpenStudy (anonymous):

= 1/2 ln | 16 + v^2 | + C

OpenStudy (anonymous):

hmmm did u do that by substitution???? i dont normally do it that way...

OpenStudy (anonymous):

technically you don't have to since it can be recognized and there are definitely other valid approaches. but i find it to be the simplest way to look at it.

OpenStudy (anonymous):

it was 1/2 ln |u|. plugged back in u = v^2 + 16

OpenStudy (anonymous):

how did we get this bit? \[1/2 \int\limits_{?}^{?} u/du\]

OpenStudy (anonymous):

sorry :( my integrations really bad x

OpenStudy (anonymous):

we had \[\int\limits_{}^{} \frac{ v dv}{ (v^{2} + 16) }\] u = v^2 + 16 ----> du = 2v dv <-----> v dv = (1/2) du = \[\int\limits_{}^{} \frac{ (1/2) du }{ u }\]

OpenStudy (anonymous):

aaaaaaaaahhhhhhhhhhhhh i seeeeee. give me one moment to assiimilate that thought plz

OpenStudy (anonymous):

what about the v on the top of the integral??? where did that go?

OpenStudy (anonymous):

v dv = 1/2 du the substitution was chosen for that to happen

OpenStudy (anonymous):

ahh ok :) and the right hand side easily integrates to 2x ???

OpenStudy (anonymous):

so

OpenStudy (anonymous):

1/2 ln (16 +v^2) = 2x +c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then

OpenStudy (anonymous):

ln(16+v^2) = 4x +c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take e of both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait i forgot to mention that x=0 when v=0, so I need to find C first

OpenStudy (anonymous):

you don't have to find it yet. usually easier to find once you have you v(x) function at the end

OpenStudy (anonymous):

oh okay i will try that now then :)

OpenStudy (anonymous):

this next part is actually a little tricky

OpenStudy (anonymous):

oh. 16+v^2 = e^4x +c

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

taking e of both sides gives: 16 + v^2 = \[e^{4x + C } = e^{4x} e^{C}\] by law of exponents

OpenStudy (anonymous):

e^C will always give a real poisitive number so we can say that e^C = C. giving Ce^(4x)

OpenStudy (anonymous):

yeah i get the first bit about law of exponents, but i dont get the second bit ...

OpenStudy (anonymous):

oh wait...

OpenStudy (anonymous):

the same way that when you multiplied both sides by 2 you got 4x + C instead of 4x + 2C. 2C is pretty redundant since C is an arbitrary constant that we don't really care how it was obtained. e^C can give any positive number in the world. same way that C can be any number in the world

OpenStudy (anonymous):

oh yeaaaahhh.... i see :)

OpenStudy (anonymous):

so

OpenStudy (anonymous):

erm, now what do we do., we have now got

OpenStudy (anonymous):

16+V^2 = Ce^(4x

OpenStudy (anonymous):

not quite sure how to find C

OpenStudy (anonymous):

(i am now your fan btw :))

OpenStudy (anonymous):

C will be the last thing you find suing the initial condition v(0) = 0. you want to isolate v. this will give you a function v(x) which is the solution to the differential equation

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

using*

OpenStudy (anonymous):

erm....speak english??? sorry no comprendo last sentence

OpenStudy (anonymous):

lol. you need to solve for v. isolate v bu itself on one side

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

i take 16 over to the other side of eqn, then sqrt to get v=.....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[v=\sqrt{Ce ^{4x} -16}\]

OpenStudy (anonymous):

yes. now you know that at x = 0. v = 0 [ v(0) = 0 ]

OpenStudy (anonymous):

this will give you C which is unique to that initial condition

OpenStudy (anonymous):

0=sqrt(C-16)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

C^(1/2) = 4

OpenStudy (anonymous):

c=16?

OpenStudy (anonymous):

ya c = 16

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ln(16+V^2) = 4x + 16 ?????

OpenStudy (anonymous):

oh actually no

OpenStudy (anonymous):

v=?

OpenStudy (anonymous):

the final equation is the solution to your differential equation yes v(x)

OpenStudy (anonymous):

\[v=\sqrt{16e ^{4x }-16}\]

OpenStudy (anonymous):

yes. it can be simplified a bit nice but it is correct

OpenStudy (anonymous):

v= 4e^(4x) -4

OpenStudy (anonymous):

urghhhh... why did you take the 4 outside the sqrt?

OpenStudy (anonymous):

oh yeah yeah yeah right sorry yeah

OpenStudy (anonymous):

so

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

v= 4 (sqrt e^(4x)) -4

OpenStudy (anonymous):

oh this is getting tricky

OpenStudy (anonymous):

erhhh... that's the answer i think??? :)

OpenStudy (anonymous):

\[v(x) = 4\sqrt{e^{4x} - 1}\]

OpenStudy (anonymous):

what i said earlier may have been misleading. deleted it

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b} = \sqrt{ab}\] \[\sqrt{16e^{4x} - 16} = \sqrt{16}\sqrt{e^{4x} - 1}\]

OpenStudy (anonymous):

ah Euler, you are a genius!!!! wow

OpenStudy (anonymous):

i finally got it.

OpenStudy (anonymous):

wow, thank u soo much. i am revising for A levels. its very tricky, but you are so very helpful :)

OpenStudy (anonymous):

i love this website too

OpenStudy (anonymous):

i'm glad i could help :) i love math and helping

OpenStudy (anonymous):

you're doing a great job helping loads of people all over the world. superman. I really appreciate it

OpenStudy (anonymous):

Thank you for your patience and everything :) <3 i will definitely come back again soon

OpenStudy (anonymous):

glad to hear that :) enjoy math

OpenStudy (anonymous):

r u a university graduate btw???

OpenStudy (anonymous):

or a professor of math or something?

OpenStudy (anonymous):

currently an undergraduate.[bachelor's then leaving]

OpenStudy (anonymous):

no. i just love math. i tutor friends

OpenStudy (anonymous):

oh wow. oh ah ok. its really good that u love math. keep spreading the love.

OpenStudy (anonymous):

hehe will do :)

OpenStudy (anonymous):

ok shall i go now...?

OpenStudy (anonymous):

yeah. i think ill give it a rest now. Thank you sooo much.

OpenStudy (anonymous):

see you . many thanks. bye xxxxxxxxx